OpenStudy (anonymous):

|2n + 4| = 8 n=2 and n=4 ?

OpenStudy (shamil98):

2n + 4 = 8 2n = 4 n = 2 2n + 4 = -8 2n = -12 n = -6

OpenStudy (anonymous):

n=2 2*2=4+4=8